OpenStudy (anonymous):

Starting at t = 0, a particle moves along a line so that its position, in metres, after t seconds, is given by s(t) = t2 − 7t + 6. Find the velocity of the particle when its position is 14

OpenStudy (campbell_st):

let s(t) = 14 and then solve the quadratic. this will give the time(s) when the particle is 14 from the origin. next find the 1st derivative, when you have that equation substitute your value(s) of t to find the velocity when the particle is 14 from the origin.

OpenStudy (anonymous):

when i solve the quadratic i get, t = 15... is that right?

OpenStudy (campbell_st):

well you have to solve 14 = t^2 - 7t + 6 or t^2 - 7t - 8 = 0

OpenStudy (anonymous):

yeah... i'm still getting 15

OpenStudy (anonymous):

At least you can factor it !!!

OpenStudy (anonymous):

lol

OpenStudy (campbell_st):

ok... do you know how to solve a quadratic..?

OpenStudy (anonymous):

well apparently not. i thought i did... :(

OpenStudy (campbell_st):

so factors of -8.. any thoughts..?

OpenStudy (anonymous):

are they not 1, 2, 4, and eight?

OpenStudy (anonymous):

oh pellet, hold on i think i got it

OpenStudy (anonymous):

nope nevermind...

OpenStudy (campbell_st):

well you need -8... not 8... and then you need to find the 2 factors that then add to - 7

OpenStudy (campbell_st):

so you can use -2 x 4, 2 x -4 , -1 x 8 or 1 x -8...

OpenStudy (anonymous):

(t - 8)(t - 1)

OpenStudy (anonymous):

i mean t + 1

OpenStudy (campbell_st):

nope... that gives t^2 - 9t + 8 = 0 its actually (t - 8)(t + 1) = 0 so only consider the solution t = 7 as time can't be negative. so the particle is 14 from the origin when t = 7 so can you find the 1st derivative...?

OpenStudy (anonymous):

2t - 7 ?

OpenStudy (campbell_st):

thats good, the 1st derivative is the rate of change in displacement... or the velocity equation. so to find the velocity when the particle is 14 from the origin... substitute t = 7 into the 1st derivative.

OpenStudy (anonymous):

then i'll get 2(7) - 7, which is 14

OpenStudy (anonymous):

but the answer is 9... unless i messed up somewhere...

OpenStudy (campbell_st):

2 x 7 - 7 isn't 14

OpenStudy (anonymous):

i meant 7 lol

OpenStudy (anonymous):

don't know what i was thinking.

OpenStudy (anonymous):

oh i was supposed to use t = 8, not 7

OpenStudy (campbell_st):

oops you need to substitute t = 8... not t = 7.... t = 8 is a solution to (t -8)(t+1) = 0 so the velocity is 2(8) - 7 = 9 units per second

OpenStudy (anonymous):

thank you